# Pattern formation and spatiotemporal chaos in relativistic degenerate plasmas

### v1

@@ V Discussion and conclusion
To elucidate the characteristics of the wave-wave interactions in degenerate plasmas, we consider three different values of the degeneracy parameter: $R_{0}=5,~{}25$, and $85$ to characterize moderate, strong and ultra-relativistic degeneracy regimes with degenerate electron number densities $n_{\rm{d0}}\sim 8\times 10^{31}~{}\rm{cm}^{-3}$, $9\times 10^{33}~{}\rm{cm}^{-3}$, and $4\times 10^{35}~{}\rm{cm}^{-3}$ respectively. In all these cases, the pattern selection restricts the number of excited harmonic modes to be smaller than $[k^{-1}]$. Also, when the modulation wave number is slightly smaller than $k_{m}$ or the length scale of excitation of harmonic modes is a bit larger than that of the most unstable mode, a few harmonic patterns coexist with the master mode. The amplitudes and widths of the solitary patterns change with time. However, the collisions and fusions among them can not occur frequently, and the spatial partial coherence of the patterns may be there, leading to the coexistence of TC and SPC states. Although the energy transfer occurs from initially excited multiple modes to a single or couple of higher harmonic modes with stronger amplitudes but reduced wavelengths, the transfer rate is not fast enough to cause STC. We have verified the existence of these ST and SPC states by analyzing the largest Lyapunov exponent spectra over the entire spatial domain and measures of the correlation function and the mutual information function over the distance (and as it increases) between any two points in space. However, as we further reduce $k$ from a starting value of $k$ $(<k_{m})$ at which the TC and SPC states coexist, there appear to be the excitation of many solitary patterns, which, as time goes on, interact themselves and fuse to form new incoherent patterns and undergo through strong electron-acoustic wave emission. The wave energy transfer occurs from initially excited many coherent modes at large scales to a single or a couple of stable higher harmonic incoherent modes with shorter wavelengths. Such a process develops faster as one approaches moderate to ultra-relativistic degeneracy regimes. The largest Lyapunov exponents are found positive in all these degenerate regimes and the vanishing of the correlations and mutual information functions over larger spatial domains.
@@
To conclude, depending on the modulational perturbation wavelength smaller or larger than that of the most unstable wave, the electromagnetic radiation spectra emanating from compact astrophysical objects and interacting with electron-acoustic perturbations sustained by multi-component degenerate plasmas that surround the compact objects may either settle into EM envelope solitons (or stable wave states) due to saturation of the modulational instability or undergo through strong electron-acoustic wave emissions causing to emerge spatiotemporal chaotic states, which may be signature of turbulence in dense astrophysical regimes [17, 16, 5].
@@ Acknowledgments
Conflict of Interest
The authors have no conflicts to disclose.
Author Contributions

### v2

@@ 5 Discussion and conclusion
To elucidate the characteristics of the wave-wave interactions in degenerate plasmas, we consider three different values of the degeneracy parameter: $R_{0}=5,~25$, and $85$ to characterize moderate, strong and ultra-relativistic degeneracy regimes with degenerate electron number densities $n_{\rm{d0}}\sim 8\times 10^{31}~\rm{cm}^{-3}$, $9\times 10^{33}~\rm{cm}^{-3}$, and $4\times 10^{35}~\rm{cm}^{-3}$ respectively. In all these cases, the pattern selection restricts the number of excited harmonic modes to be smaller than $k^{-1}]$. Also, when the modulation wave number is slightly smaller than $k_{m}$ or the length scale of excitation of harmonic modes is a bit larger than that of the most unstable mode, a few harmonic patterns coexist with the master mode. The amplitudes and widths of the solitary patterns change with time. However, the collisions and fusions among them can not occur frequently, and the spatial partial coherence of the patterns may be there, leading to the coexistence of TC and SPC states. Although the energy transfer occurs from initially excited multiple modes to a single or couple of higher harmonic modes with stronger amplitudes but reduced wavelengths, the transfer rate is not fast enough to cause STC. We have verified the existence of these ST and SPC states by analyzing the largest Lyapunov exponent spectra over the entire spatial domain and measures of the correlation function and the mutual information function over the distance (and as it increases) between any two points in space. However, as we further reduce $k$ from a starting value of $k$ $(<k_{m})$ at which the TC and SPC states coexist, there appear to be the excitation of many solitary patterns, which, as time goes on, interact themselves and fuse to form new incoherent patterns and undergo through strong electron-acoustic wave emission. The wave energy transfer occurs from initially excited many coherent modes at large scales to a single or a couple of stable higher harmonic incoherent modes with shorter wavelengths. Such a process develops faster as one approaches moderate to ultra-relativistic degeneracy regimes. The largest Lyapunov exponents are found positive in all these degenerate regimes and the vanishing of the correlations and mutual information functions over larger spatial domains.
An important finding of this work is that, in ultra-relativistic degenerate plasmas $(R_{0}\gg 1)$, the modulational instability growth rate does not always saturate within the domain $0<k<1$ (Note that since $k$ is normalized by the inverse of the plasma skin depth, the values with $k>$ are inadmissible, otherwise, the plasma collective behaviors may disappear) into stable solitons as suggested in earlier studies [shatashvili2020nonlinear], but may evolve into incoherent spatiotemporal chaos by the influence of electron-acoustic wave emission, thereby establishing a new dynamical regime beyond the soliton regime.
We mention that the spectral broadening and rapid variability in EM wave emissions from compact astrophysical objects are typically related to nonlinear wave-particle interactions. These phenomena can be associated with chaotic states and broadband power spectra. Nonlinear wave-particle interactions can cause frequency broadening and chaotic variations in radiation spectra and their characteristics. Although nonlinear wave-wave interactions can also contribute to spectral changes, especially rapid variability, these are usually the result of the superposition of different waves and may not be related to the interactions studied here. Another important phenomenon, EM pulse fragmentation, is also primarily related to wave-particle interactions. In compact objects, extreme conditions lead to fragmentation of matter, which is influenced by EM waves or particles. Processes like scattering and absorption, where energy transfers between waves and particles, can drive fragmentation.
@@
To conclude, depending on the modulational perturbation wavelength smaller or larger than that of the most unstable wave, the electromagnetic radiation spectra emanating from compact astrophysical objects and interacting with electron-acoustic perturbations sustained by multi-component degenerate plasmas that surround the compact objects may either settle into EM envelope solitons (or stable wave states) due to saturation of the modulational instability, especially in the regime of weak relativistic degeneracy or undergo through strong electron-acoustic wave emissions causing to emerge spatiotemporal chaotic states, especially in the regime of ultra-relativistic degeneracy, which may be signature of turbulence in dense astrophysical regimes [banerjee2010spatiotemporal, misra2009pattern, chian1999]. Zakharov-like models are limited to studying the generation of envelope solitons through MI, wave collapse, and wave turbulence through wave-wave interaction. However, phenomena such as the generation of dromion solutions or vortex formation, as well as those related to wave-particle interactions, cannot be studied using Zakharov-like equations. Building on these limitations, our results show that STC can arise as a natural state in relativistic degenerate plasmas. These chaotic states may explain the origin of turbulence and fluctuations in radiation in dense astrophysical environments. Future studies with two- or three-dimensional models and comparisons with observational data will be essential to confirm these results and understand their relevance in real astrophysical systems.
@@ Acknowledgments
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
CRediT authorship contribution statement
@@
Appendix A Pseudocode to reproduce numerical integration
We recast Eqs. (4) and (5) as
$\begin{split}\frac{\partial A}{\partial t}=-v_{g}\frac{\partial A}{\partial x}+\frac{i}{2}\left(B_{0}\frac{\partial^{2}A}{\partial x^{2}}+B_{1}NA+B_{2}|A|^{2}A\right)\equiv F_{1},\end{split}$ (17)
$\frac{\partial N}{\partial t}=U\equiv F_{2},$ (18)
$\frac{\partial U}{\partial t}=\frac{\partial^{2}N}{\partial x^{2}}-3b_{3}\frac{\partial^{2}|A|^{2}}{\partial x^{2}}\equiv F_{3}.$ (19)
In our numerical algorithm, we discretize time as $t=t_{k}=k\,\Delta t,\quad k=0,\ldots,N_{t},$ and space as $x=x_{j}=j\,\Delta x,\quad j=0,\ldots,N_{x}-1.$ We denote the solution as $u(x_{j},t_{k})\equiv u^{k}_{j}$, and use the periodic boundary conditions: $A(L,t)=A(0,t),$ $N(L,t)=N(0,t),$ and $U(L,t)=U(0,t),$ or, for the discretized solution, $A^{k}_{N_{x}-1}=A^{k}_{0}$, $N^{k}_{N_{x}-1}=N^{k}_{0}$, and $U^{k}_{N_{x}-1}=U^{k}_{0}$. The functions $F_{1}$, $F_{2}$ and $F_{3}$ contain discretized variables where we use the central difference approximations for the spatial derivatives, i.e.,
$\frac{\partial A}{\partial x}\approx\frac{A_{j+1}-A_{j-1}}{2\Delta x},\quad\frac{\partial^{2}A}{\partial x^{2}}\approx\frac{A_{j+1}-2A_{j}+A_{j-1}}{\Delta x^{2}},$
for $j=1,2,\cdots N_{x}-2$. At the boundaries ($j=0$ and $j=N_{x}-1$) the spatial derivatives are defined as
$\left.\frac{\partial A}{\partial x}\right|_{x_{0}}\approx\frac{A_{1}-A_{N_{x}-1}}{2\Delta x},\quad\left.\frac{\partial^{2}A}{\partial x^{2}}\right|_{x_{0}}\approx\frac{A_{1}-2A_{0}+A_{N_{x}-1}}{\Delta x^{2}},$
$\left.\frac{\partial A}{\partial x}\right|_{x_{N_{x}-1}}\approx\frac{A_{0}-A_{N_{x}-2}}{2\Delta x},$
$\left.\frac{\partial^{2}A}{\partial x^{2}}\right|_{x_{N_{x}-1}}\approx\frac{A_{0}-2A_{N_{x}-1}+A_{N_{x}-2}}{\Delta x^{2}},$
where $\Delta x$ is the spatial step, and similar expressions for $N$ and $U$.
Next, we define the triplet, $\left(F_{1},F_{2},F_{3}\right)$ as the vector-valued function $\mathbf{F}\left(A,N,U,t\right)$, i.e.,
$\left(F_{1},F_{2},F_{3}\right)=\mathbf{F}\left(A,N,U,t\right).$ (20)
Then the fourth-order Runge–Kutta scheme reads
$(k_{1A},k_{1N},k_{1U})$ $\leftarrow\mathbf{F}(A,N,U,t),$
$(k_{2A},k_{2N},k_{2U})$ $\leftarrow\mathbf{F}\!\Big(A+\tfrac{\Delta t}{2}k_{1A},\;N+\tfrac{\Delta t}{2}k_{1N},$
$\hskip 48.36958ptU+\tfrac{\Delta t}{2}k_{1U},\;t+\tfrac{\Delta t}{2}\Big),$
$(k_{3A},k_{3N},k_{3U})$ $\leftarrow\mathbf{F}\!\Big(A+\tfrac{\Delta t}{2}k_{2A},\;N+\tfrac{\Delta t}{2}k_{2N},$
$\hskip 48.36958ptU+\tfrac{\Delta t}{2}k_{2U},\;t+\tfrac{\Delta t}{2}\Big),$
$(k_{4A},k_{4N},k_{4U})$ $\leftarrow\mathbf{F}\!\Big(A+\Delta t\,k_{3A},\;N+\Delta t\,k_{3N},$
$\hskip 48.36958ptU+\Delta t\,k_{3U},\;t+\Delta t\Big),$
Thus, we obtain the solutions for $A$, NN, and $U$ at time $t+\Delta t$ as
$A$ $\leftarrow A+\tfrac{\Delta t}{6}\left(k_{1A}+2k_{2A}+2k_{3A}+k_{4A}\right),$ (21)
$N$ $\leftarrow N+\tfrac{\Delta t}{6}\left(k_{1N}+2k_{2N}+2k_{3N}+k_{4N}\right),$ (22)
$U$ $\leftarrow U+\tfrac{\Delta t}{6}\left(k_{1U}+2k_{2U}+2k_{3U}+k_{4U}\right).$ (23)
Typically, an explicit time integration scheme exhibits stability when the time step, $\Delta t$ is chosen sufficiently small such that the condition, $\Delta t<C\,(\Delta x)^{p}$ holds for some positive constant $C$. Here, the exponent, p(>0)p~(>0) corresponds to the order of the highest-order spatial derivative in the evolution equation [Cross_Greenside_2009]. For the present model, $p=2$ and $C={\min}_{j}\left\{1/2D_{j}\right\}$ with $D_{j}$ denoting the moduli of the coefficients of the highest derivatives. In Eqs. (17)–(19), there are three second-order spatial derivatives, namely, $\partial^{2}A/\partial x^{2}$, $\partial^{2}N/\partial x^{2}$, and $\partial^{2}|A|^{2}/\partial x^{2}$ with coefficients $iB_{0}/2$, $1$, and $3b_{3}$ respectively. Thus, the stability condition reduces to
$\begin{split}\Delta t<&\min\left\{\frac{\Delta x^{2}}{B_{0}},\frac{\Delta x^{2}}{2},\frac{\Delta x^{2}}{6b_{3}}\right\}\\ &=\Delta x^{2}\min\left\{\frac{1}{B_{0}},\frac{1}{2},\frac{1}{6b_{3}}\right\}\equiv\tau_{\min}.\end{split}$ (24)
In our simulation, we chose the time step as
$\Delta t<\min\left\{10^{-4},\tau_{\min}\right\}.$ (25)
Also, since we considered the simulation box size as $L_{x}=2\pi/k$ and the number of grid points as $2048$, the spatial step size is $\Delta x=L_{x}/2048$.
$R_{0}$ $k$ $\Delta x$ $\Delta t$
55 $0.15$ $0.0205$ $2\times 10^{-5}$
$0.11$ $0.0279$ $3\times 10^{-5}$
$0.043$ $0.0713$ $10^{-4}$
$25$ $0.21$ $0.0146$ $5\times 10^{-5}$
$0.11$ $0.0279$ $10^{-4}$
$0.047$ $0.0653$ $10^{-4}$
$85$ $0.29$ $0.0106$ $5\times 10^{-5}$
$0.11$ $0.0279$ $10^{-4}$
$0.047$ $0.0653$ $10^{-4}$
Table 3: Spatial $(\Delta x)$ and time $(\Delta t)$ steps for each simulation using Eqs. (25).